259 views
 owned this note
--- title: Minutes of the PS & SPS User Meeting --- # PS-SPS Users Meeting for Week 18 held on 30 April 2025 Coordinator for week 17 & 18 : Paolo Martinengo :::success * Current **PS / SPS User schedule v1.1.1** Problem with updating web page solved. Schedule to be put on web page after the meeting. * **Injector schedule** (v1.1) for 2025 is available here: https://edms.cern.ch/document/3057281/latest ::: :::warning **EuroLabs :** We will soon send out the call for 2025. New day rates (tbc). ::: ## News from the PS & SPS Physics Coordinator M.R. Jäkel, P. Martinengo :::info **Proton Run 2025** * Physics Start p+ EA & nTOF Wednesday 19.03.2025 * Physics start Isolde Friday 28.03.2025 * SPS NA Physics start EHN1 Monday 14.4.2025 * SPS NA Other Physics start Thursday 17.4.2025 * LHC start delayed by ≈ 1 week ::: :::success **Newly vacant beam slots :** * PS T10 - W16, W21-22 * SPS H2 - W16-17 * SPS H4 - W40 (beam restrictions apply) * SPS H6 - W33 * SPS H8 - W23 Please contact [SPS.Physics.Coordinator@cern.ch](mailto:SPS.Physics.Coordinator@cern.ch) if you want to apply for any of theses slots. ::: Reminder: Beam time exceeding the limits of 2 weeks PS beam time and 1 week SPS beam time per year need the approval of one of these CERN committees: SPSC, LHCC, DRDC, INTC, RB or IEFC. Consider joining a DRD collaboration, if you require more beam time. **Target intensities:** See below or via [ASM](https://asm.cern.ch/experimental-area/experiments) (accessible from within the CERN network). | T2 | T4 | T10 | T6 | Date | Week | | --- | --- | --- | --- | ------ | ---- | | 100 | 45 | 21 | 50 | 30.04. | 18 | | 100 | 45 | 21 | 50 | 07.05. | 19 | | 100 | 38 | 21 | 50 | 14.05. | 20 | | 100 | 38 | 21 | 50 | 21.05 | 21 | | 100 | 38 | 21 | 150 | 28.05. | 22 | ![](https://codimd.web.cern.ch/uploads/upload_9294cc92f03f0dc6edd2e31eac2d132e.png) ### News from the Facilities Operations Meeting (FOM) :::danger SPS : Problems establishing channeling using the crystal (TECS.21602) which should shadow the ZS. Maximum intensity to NA limited. ::: **Upcoming SPS MDs next weeks :** https://be-dep-ea.web.cern.ch/content/md-planning-north-area Dedicated SPS MDs - if not anounced differently - are taking place from 8:00 to 18:00 --- ## PS Machine Report (Ewen Hamish maclean) ![](https://codimd.web.cern.ch/uploads/upload_d6e46807b019d05e25e9c4e9afd0e9d6.png) Decent availability over last 7 days: ~95% Main issues in PS: - fault in C10-56 led to repeated trips throughout Monday. Ran on spare 10MHz Cavity overnight then scheduled access on Tuesday morning to coincide with planned interventions on TN (beam stop was required at this point anyway to avoid risk of temporarily loosing control of CCC consoles) - access done 11.30-12.30 on Tuesday. No major cavity issues since. - Following TN interventions on Tuesday took some time for Beam to be recovered in PSB (issue with reset of BIS). - some issues with bunch splitting, but affecting LHC/SPS-MD beams Various optimizations on operational beams during the week. - barrier bucket adjustments for MTE - improvements to losses at injection on EAST cycles by optimizing tune - TOF re-steering ## SPS Machine Report (Pablo Arrutia) Availability (Thu - Wed): ~91% Main SPS faults - Mon: Injection kicker (MKP) generator fault (~2.5h) - Tue, yesterday: Vacuum degradation on Tbid T2 after intensity increase (under investigation) This week - AWAKE run 1 - LHC filling - NA physics: - Mon: RF and OP experts worked on spill optimisation to remove [bump](https://indico.cern.ch/event/1540836/contributions/6484453/subcontributions/546117/attachments/3056032/5403042/SPSUserMeeting_24042025.pdf). **Need feedback from users:** ran with new settings Mon 12:30pm to Wed 1:30am, then reverted. - Tues: Doubled extracted intensity and adjusted sharing - Wed, today: dedicated MD -> No beam for NA until 18h - Upcoming (may impact NA): alignment of SPS operational crystal (TECS) - might request ~1-2h at reduced intensity - motor is now repaired - important for long term operation to reduce beam loss at extraction Next week: - HiRadMat - LHC scrubbing - NA physics ## Safety (James Devine) Thanks for everyone who submitted their ISIEC requests well ahead of time, and for letting us know when your schedules have changed. We are still very busy with experiments outside of test beams, if you haven't heard from us for a confirmation please do send a reminder and we will get back to you as soon as possible. As a reminder, if you are running for longer than 3 weeks we can give you a longer safety clearance with a slightly different procedure, please contact us for details. Our team email address is [EP TH Safety](mailto:EP-TH.Safety@cern.ch) Finally, with the warm weather approaching a reminder that for work in beam line areas - where safety shoes are required, shorts are not permitted. Depending on your activity, t-shirts are permitted but your shoulders must be covered (i.e. no singlets). Link to safety requirements on clothing for heatwaves: [EDMS 2193556](https://edms.cern.ch/ui/file/2193556/1/HSE_memorandum_Habit_de_travail_EDMS_2193556.pdf) ## TT42 Main: n_TOF (Alice Manna, Paolo M Milazzo) In general a very smooth data taking. Today (30.04) in EAR1 end of the 97,98Mo(n, gamma) data taking campaign and start of the 87Sr(n, gamma) measurement. Same experimental set-up (C6D6 liquid scintillators) will be used. Calibration procedures needed and a stop of approximately 6 hours is estimated. Data taking will continue as before in other experimental areas: • EAR2: study of the 124Sn(n, gamma) reaction • NEAR: Study of radiation hardness for a diamond detector ![](https://codimd.web.cern.ch/uploads/upload_357ea68bb9c25f6443836c9e4e861dab.png) Protons received per Week ## East Area Beam Status (B. Rae) On call phone number: **67500** General : Smooth Operation T9: Good operation. T10: Good operation. In the end, we will not go for the longer spill duration from today as ALICE TIMING agreed with ALICE ITS3 to stay a little longer and swap on Friday / Saturday. T11: No user. ## East Area Users Tour de Table ### T8 Main: IRRAD/CHARM (F. Ravotti / S. Fiore) Overall good week, 2.4E16 protons delivered to T8 (109% of the target) and CHARM reference TID exceded. Also very good beam centering (95-96%). ![](https://codimd.web.cern.ch/uploads/upload_b8d646c8729ea94db9e25a638049f249.png) Beam stopped on Tuesday to allow for a longer cool-down to install a special experiment at IRRAD on Wednesday. IRRAD: new LHCb installation today, all the other expriments and samples continue. CHARM: remove setup by ALICE, installed permanent magnet samples, changed setups by BE-CEM. We migh have a special access next week to remove dosimetry samples (to be scheduled) and regular access next Wednesday. ### T9 Main: WCTE (Laurence Cook) Overall good operation this week: * Continued taking pion data at +260-280 MeV for pion scattering in water measurement * Took extended runs at - 250 MeV, and -300 MeV and -340 MeV using new trigger configuration reported last week for measuring Li9 decays produced by spallation of the Oxygen nucleus in water by pion (background to diffuse supernova neutrino detection). * Changed configuration to take tagged-gamma data, using a neodymium magnet to deflect the beam onto hodoscope and allowing us to look for gamma 100 MeV - 1 GeV produced via Bremsstrahlung in our beam monitors with the water cherenkov detector, * We completed 4 10 hour runs at 500 MeV, 650 MeV, 800 MeV and 1 GeV over the weekend * Monday evening we had 7 hours without the beam to allow for the installation of our new triggering mode that distributes the trigger signal directly to mPMTs in the tank allowing for taking data at higher rates and for the readout to be less affected by high energy backgrounds from other beamlines. * Commissioning of this new readout mode started yesterday evening taking data at 780 MeV - more work ongoing checking data quality and testing higher data taking rates. Preliminary event display of a high energy muon entering our detector ![](https://codimd.web.cern.ch/uploads/upload_1c2f92ec258df4d014afbb595c1095cc.png) Plans for this week: * Continued testing and data taking with the hardware trigger * Continue taking lower energy data with 9Li trigger while work to ensure data quality with new external trigger mode * Start taking data at higher rates for muon scattering analyses with the new external trigger ### T10 Outgoing Main: ALICE_TIMING (Manuel Colocci) - Took data with SiPMs (FBK prototypes) with a scope in the first days, and since Tuesday with a digital readout (picoTDC) at -20C (and LGADs reference planes). - Setup unistalled this morning. ### T10 Main: ALICE_ITS3 (Paolo Martinengo) intervention on the DESY table ongoing (till noon), will be ready for beam just after (and pending safety inspection) Long spill not required, keep present setting ### T10 Incoming Main: SHIP LS-SBT (Please insert your name) ## North Area Beam Status (B. Rae) On call phone number: **67500** General: T4 target head to be changed today. H2: No user scheduled, LHCb starts today. H4: Mostly smooth operation, NA64e beamline modifications ongoing today. PPE144 door showed "unsafe" state despite being closed. Piquet intervened and it was fixed. MBPL022544 was found with a valve closed. EHN1 stopped for 45 mins and access in the EHN1 galleries was organized on 25.04 by BE-EA and TE-MSC. The problem was resolved. H6: Good operation. H8: Good operation. M2: Electron beam checked. Final checks to be done with the user on 30th April. P42/K12: No issues. Changing the sharing led to better transmission in P42 (from T4 to T10), so intensity for NA62 needed to be reduced yesterday evening. ## TT66 Main: HIRADMAT (A. Goillot) * DPA experiment installed in TNC, latest issues being resolved. * Beam starting on Monday, optics tested. ## TT41 Main: AWAKE (M. Bergamaschi) In general last week limited number of extractions due to competing commissioning of injections in LHC. Detailed program: * Wednesday: E-beam in plasma studies * Thursday: Measurements with SPS RF expertes for 1hr+ E-beam in plasma studies * Friday: E-beam in plasma + few extraction * Saturday: No extractions to AWAKE to competing LHC injection tests * Sunday: proton beam studies + stop due to LINAC4 issue * Monday: physics program with plasma * Tuesday: physics program with plasma, beam down for 2 hours for network reboot ## TELMAX Incoming Main: PAX (Michael Roosa) PAX setup is installed and under vacuum in TELMAX. Prototype TES detector succesfully arrived and deployed with NIST team. Safety inspection passed, beam permit obtained. Currently tuning and testing new antiproton extraction mode (MMBE--Multi Mini-Bunch Extraction) with AD operators into the target chamber. Noise chasing in the TES ongoing. Physics beamtime planned from May 5-19th for antiprotonic atom x-ray spectroscopy with solid targets. Here is a CAD drawing of our test beam setup. The antiproton beam arrives from the floor and impinges on the solid targets. ![](https://codimd.web.cern.ch/uploads/upload_040afee11d3b0f5571a4a69f9f4e64cf.png) Here are some photos of PAX setup in TELMAX. (Left)The prototype TES microsnout, behind its closed x-ray shield, being installed in the cryostat. (Middle) The cryostat closed with the TES inside, and moved into its final position outside the Ti window close to the interaction region. (Right) The team for test beam in the TELMAX zone with the full experimental setup. ![](https://codimd.web.cern.ch/uploads/upload_5bac9bd120b31bf244fcf0e13785df45.png) ## North Area Users Tour de Table ### H2 Main: LHCB (D. Manuzzi) Goal: characterisation of timing detector based on MCP technology - No particular news compared to last week report. - Installation is ongoing. All detectors (DWC, scintillators, vacuum chamber with prototype) on the DESY taable. We expect to be ready for the safety visit today at 15:00, as booked. - Plan to start taking data as soon as possible. - Beam file: we asked for 50 GeV electrons and muons (whatever energy). Many thanks for the support. ### H4 Main: NA64e (L. Molina Bueno) We passed our safety inspection last friday and we continue with commissioning and calibration of the detectors since then. This year we are using new waveboards to read out all calorimeters. Today the beam material along the line and the extra vaccum will be installed and the water leak observed from the third magnet will also be fixed today. Since yesterday afternoon we have also 100 units on T2 target. We are going to try to complete calibrations and start data taking in the following days. We would like to thank all the help from BE-EA department specially Nikos and Sylvain. ### H6 Outgoing Main: ATLAS_HGTD (Stefano Manzoni) 16 days of data taking. After the initial issue with the chiller, quickly resolved by Andre (many thanks!), everything ran smoothly. The main achievement of this test beam was the first parallel test of three HGTD modules. Data analysis is currently ongoing. ![](https://codimd.web.cern.ch/uploads/upload_b4da1cfad9e06f92abbd8beaca4d2504.jpeg) Many thanks to AFP and AIDAInnova for their collaboration and support with beam setup and access management over the past weeks. ### H6 Main: ATLAS_ITK_PIXEL (Andre Rummler) We are starting to set up during this meeting. Handover from HGTD done on time, thanks a lot to everybody. We will set up for the first time ITk pixel linear triplet modules for which we need to try out new holders. Swap of chillers for the cold box to be done. Safety visit targeted for 1730 today. ### H6 Incoming Main: ATLAS_BCM_PRIME (Please insert your name) ### H6 Incoming Main: CMS MTD ETL ETRO Federico Siviero First week of a 2-weeks long data taking. Aim of the test beam is to prepare for the ETROC PRR. We will be testing 8 new ETROC2.01/HPK test boards. We plan to test them at CERN with extensive thermal cycles (-50C to 50C, if feasible, beyond -50C later). We will also study the timing performance. There might be a chance we may have few ETROC2.03/HPK test boards ready at last minute. We will use the AIDA telescope (during the 2nd week of data taking) and the associated infrastructure: cool box, chiller and piezo-stages, commissioned in April. We requested a "XSCA" (yellow table) where to sit our setup for the first week. During the second week, the setup will move to the AIDA telescope. During the first week, the AIDA telescope will be used by the other ETL team running parasitic. We will need dry air. ### H6 Outgoing Parallel: ATLAS_AFP_TOF (Tomáš Komárek) - working on PicoTDC calibration - staying in as parasitic for another week following agreement over email (no incoming users in the H6B/PPE156 zone) ![](https://codimd.web.cern.ch/uploads/upload_948422d78e2536c8e01d18338f2c6f0d.png =x220) ![](https://codimd.web.cern.ch/uploads/upload_d188b53d2e1ab012e37b4f0bdaadd046.png =x220) ### H8 Outgoing Main: LHCB (Please insert your name) ### H8 Main: SHIP CAL (Please insert your name) ### K12-P42 Main: NA62 (Roberto Piandani) No major issues during the week. Few periods of bad quality beam but from the machine they worked to improve the situation specially on Monday, after the tning was nice and on Tuesday, after the tuning we had a nice structure but with some "bump" at the extremes of the extraction. Possible problems with fake timing on Tuesday (28/04) when the MD4 user was in the SC ### M2 Main: MUONE (Aldo Arena) The 29th of April we installed the Desy table and the electromagnetic calorimeter. This morning we have just had the safety clearence for partial setup and this afternoon we plan to use 40 GeV electron beam to test and calibrate the ECAL while we will also start synchronization between tracker stations (2 installed so far). The beam will be used until the 5th of May when we will finalize the setup and installation of the insulating tent. ![](https://codimd.web.cern.ch/uploads/upload_cb1e76e48d0fcc6576aa5c02e2c03f80.jpg) ![](https://codimd.web.cern.ch/uploads/upload_4017c09e7aae6eb41e5af4d9c7355dfc.jpg) ![](https://codimd.web.cern.ch/uploads/upload_576d691b7976540d528dc10c80b02df4.jpg) ### H6 Incoming Parasitic: CMS MTD ETL ETRO Federico Siviero Running parasitic to the other ETL group. Aim of the first part of the week is to test small LGAD prototypes from the latest ETL sensor productions (FBK, HPK), preparing for the sensor pre-series. We will be measuring collected charge, timing resolution, efficiency. In the second part of the week, we will change setup and start measuring ETL modules to characterize system level performance. Goal is to confirm recent results obtained at DESY and practice ahead of the summer test beams. We will use the AIDA telescope (both setups), placed upstream wrt the other ETL team. We have our own cold box (can fit on top of the PI stage in the AIDA telescope through a 3d-printed plate) and chiller. We will need dry air and the PI-stage that usually sits between the planes of the tracker. ### H8 Parasitic: STRAW_TRACKER_RD (Please insert your name) ### M2 Parasitic: AMBER (Please insert your name) --- ## AoB --- _Minutes by the respective speakers, edited by M. Jäkel, M. Malkin, and M. Schwinzerl_